If there are 2 girls and 3 boys, how many ways are there to arrange these 5 people such that 2 girls will not be next to each other?
Answer by ewatrrr(24785)

2 girls and 3 boys
All ways of arranging: 5!
ways two girls 'would' be together: 2*4!
Ways Not together: 5! - 2*4! = 120-48 = 72 ways